Find the area between = -2 & 3

What ia the total area tocthe left of z = 2.18?
1
Expert's answer
2017-03-28T11:35:07-0400
P(-2<Z<3)=P(Z<3)-P(Z<-2)=P(Z<3)-1+P(Z<2)=
=0.9987-1+0.9772=0.9759.


P(Z<2.18)=0.9854.